OpenStudy (anonymous):

Let f(x)=3x and g(x)=x^2-5. What is (g*f)(-2)?

OpenStudy (anonymous):

Please help! Will give medal to best answer. The * sign between (g*f) is an open circle.

OpenStudy (anonymous):

@Coolsector @satellite73 Please show steps!

OpenStudy (anonymous):

@mathstudent55

OpenStudy (jdoe0001):

\(\bf (g\times f)(-2)\implies g(-2)\times f(-2)\)

OpenStudy (anonymous):

(g*f) (-2) = g(f(-2)) = g(-6) = 36-5 = 31

OpenStudy (anonymous):

Are these all the steps? @Coolsector @jdoe0001

OpenStudy (anonymous):

what jadoe wrote is not correct